# Multi-Agent Collaboration in Incident Response with Large Language Models

### v1

@@ 3 Incident Response Game
Backdoors & Breaches (B&B) is a tabletop card game designed to simulate the complexities of cybersecurity incident response in a collaborative and engaging manner. Developed by Black Hills Information Security and Active Countermeasures (2020), the game offers a hands-on approach to learning key strategies for identifying, containing, and mitigating cyber threats. By combining structured gameplay with elements of randomness, B&B immerses players in realistic decision-making processes, fostering critical thinking and teamwork.
@@ 3.2 Card Types
A complete list of B&B cards can be found in Appendix A.
@@ 4.1 Study Design
This study aims to evaluate the effectiveness of LLM-based multi-agent collaboration in identifying hidden attack cards within simulated incident response scenarios. The experiment simulates realistic cybersecurity incident response dynamics using predefined agent roles, including an incident captain and multiple defenders with their system messages in Appendix B and C. The incident captain agent is responsible for setting up scenarios with hidden attack cards representing various breach stages, introducing procedure cards, managing inject events, and ensuring adherence to game rules. To encourage team decision-making and collaboration, the captain facilitates gameplay without revealing correct solutions.
@@ 5.1 Experimental Setup
We evaluated the performance of LLM-based multi-agent collaboration in incident response by simulating six distinct team structures within the B&B game using the AutoGen framework (Wu et al. 2023). Each agent, powered by GPT-4o (Achiam et al. 2023) and configured with a temperature of 1, is assigned a predefined role with tailored responsibilities. The team structures included five defenders, varying in leadership, expertise, and collaboration styles: homogeneous centralized structure: 1 team leader and 4 team members; heterogeneous centralized structure: 1 team leader, 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, and 1 deception and containment expert; homogeneous decentralized structure: 5 team members; heterogeneous decentralized structure: 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, 1 deception and containment expert, and 1 incident response expert; homogeneous hybrid structure: 3 experts and 2 beginners; heterogeneous hybrid structure: 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, and 2 beginners. We conducted 20 simulations for each team structure.
@@ 5.3 Case Studies
We analyzed 12 failure cases across different team structures to identify key challenges that impeded effective incident response. Homogeneous centralized teams often over-relied on standard procedures, failing to adapt to subtle breach indicators. Heterogeneous centralized teams struggled with prioritization, neglecting critical network-level indicators or delaying behavior-based analytics. Decentralized structures, both homogeneous and heterogeneous, frequently relied excessively on high-modifier procedures while neglecting areas like behavior analytics. Hybrid teams revealed the importance of aligning diverse expertise, with failures stemming from misaligned priorities or underutilized procedures like memory and log analysis. These findings highlight the need for adaptive and context-aware procedure selection to optimize incident response outcomes across team structures.
@@ Acknowledgments
Appendix A Backdoors & Breaches Cards
@@
Appendix B Sequence of Game
@@
Appendix C System Messages for Agents
@@
C.1 Incident Captain Agent
@@
C.2 Defender Agents

### v2

@@ 3 Incident Response Game
Backdoors & Breaches (B&B) is a tabletop card game designed to simulate the complexities of cybersecurity incident response in a collaborative and engaging manner. Developed by Black Hills Information Security and Active Countermeasures (2020), the game offers a hands-on approach to learning key strategies for identifying, containing, and mitigating cyber threats. By combining structured gameplay with elements of randomness, B&B immerses players in realistic decision-making processes, fostering critical thinking and teamwork. A detailed discussion on the relation between B&B and real-world cybersecurity incident response is provided in Appendix A.
@@ 3.2 Card Types
A complete list of B&B cards can be found in Appendix B.
@@ 4.1 Study Design
This study aims to evaluate the effectiveness of LLM-based multi-agent collaboration in identifying hidden attack cards within simulated incident response scenarios. The experiment simulates realistic cybersecurity incident response dynamics using predefined agent roles, including an incident captain and multiple defenders with their system messages in Appendix C and D. The incident captain agent is responsible for setting up scenarios with hidden attack cards representing various breach stages, introducing procedure cards, managing inject events, and ensuring adherence to game rules. To encourage team decision-making and collaboration, the captain facilitates gameplay without revealing correct solutions.
@@ 5.1 Experimental Setup
We evaluate the performance of LLM-based multi-agent collaboration in incident response by simulating six distinct team structures within the B&B game using the AutoGen framework (Wu et al. 2023). Each agent, powered by GPT-4o (Achiam et al. 2023) and configured with a temperature of 1, is assigned a predefined role with tailored responsibilities. The team structures include five defenders, varying in leadership, expertise, and collaboration styles: homogeneous centralized structure: 1 team leader and 4 team members; heterogeneous centralized structure: 1 team leader, 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, and 1 deception and containment expert; homogeneous decentralized structure: 5 team members; heterogeneous decentralized structure: 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, 1 deception and containment expert, and 1 incident response expert; homogeneous hybrid structure: 3 experts and 2 beginners; heterogeneous hybrid structure: 1 endpoint security expert, 1 network traffic analysis expert, 1 log and behavioral analysis expert, and 2 beginners. We conduct 20 simulations for each team structure, with one example of a game simulation provided in Appendix E for the homogeneous centralized team structure.
@@ 5.3 Case Studies
We analyzed 12 failure cases across different team structures to identify key challenges that impeded effective incident response, with detailed analyses provided in Appendix F. Homogeneous centralized teams often over-relied on standard procedures, failing to adapt to subtle breach indicators. Heterogeneous centralized teams struggled with prioritization, neglecting critical network-level indicators or delaying behavior-based analytics. Decentralized structures, both homogeneous and heterogeneous, frequently relied excessively on high-modifier procedures while neglecting areas like behavior analytics. Hybrid teams revealed the importance of aligning diverse expertise, with failures stemming from misaligned priorities or underutilized procedures like memory and log analysis. These findings highlight the need for adaptive and context-aware procedure selection to optimize incident response outcomes across team structures.
@@ Acknowledgments
Appendix A Relation to Real Incident Response
The Backdoors & Breaches (B&B) framework provides a structured and engaging simulation of real-world cybersecurity incident response, emphasizing critical skills such as strategic planning, team collaboration, and adaptive decision-making under uncertainty. By mirroring key stages of a cyberattack—initial compromise, pivot and escalate, command and control, and persistence—the game aligns closely with the dynamics of real-world incident response, where teams must detect threats, prioritize actions, and address unexpected challenges. Within B&B, the incident captain, analogous to a cybersecurity team lead, crafts scenarios that replicate the complexity and unpredictability of real breaches, while defenders collaboratively apply investigative procedures and respond to inject events representing disruptions or resource constraints. However, it is important to acknowledge the limitations of the framework. The game abstracts critical aspects of real incidents, such as the scale and complexity of infrastructures, the iterative nature of adversarial attacks, and the need for cross-departmental communication during live breaches. Additionally, while randomness through dice rolls and inject cards captures some unpredictability, it cannot fully emulate the evolving tactics of real attackers. Despite these limitations, B&B remains a valuable tool for practicing the core principles of incident response and serves as a foundation for exploring large language model agent collaboration in realistic cybersecurity scenarios.
Appendix B Backdoors & Breaches Cards
@@
Appendix C Sequence of Game
@@
Appendix D System Messages for Agents
@@
D.1 Incident Captain Agent
@@
The hidden attack cards for this game scenario are as follows:
[Incident Cards]
The available procedure cards are divided into Established and Other Procedures:
[Procedure Cards]
D.2 Defender Agents
@@
Appendix E Game Example
This appendix presents the real setup and turn-by-turn trajectory of one game simulation of the homogeneous centralized team structure.
E.1 Game Setup
  • Defenders: team leader defender, and team member defender 1-4.
  • Incident cards: External Cloud Access, Credential Stuffing, HTTP as Exfil, and Application Shimming.
  • Established procedure cards: Server Analysis, Endpoint Analysis, Crisis Management, and Isolation.
  • Other procedure cards: Security Information and Event Management (SIEM) Log Analysis, Firewall Log Review, Network Threat Hunting - Zeek/RITA Analysis, Cyber Deception, Endpoint Security Protection Analysis, User and Entity Behavior Analytics (UEBA), and Memory Analysis.
E.2 Game Trajectory
Table 2 presents a comprehensive breakdown of the turn-by-turn trajectory of the game, detailing the procedures used, dice rolls, outcomes, and other significant events that occurred during gameplay.
Table 2: Turn-by-turn game trajectory with the homogeneous centralized team structure.
Turn Procedure Dice Roll Modifier Success Incident Revealed Inject Event
1 Endpoint Analysis 4 +3 No - -
2 Server Analysis 16 +3 Yes Application Shimming -
3 SIEM Log Analysis 1 +0 No - Lead Handler Has a Baby
4 Isolation 14 +3 Yes HTTP as Exfil -
5 UEBA 20 +0 Yes Credential Stuffing Honeypots Deployed
6 Memory Analysis 15 +0 Yes External Cloud Access -
E.3 Game Summary
  • Total turns played: 6.
  • Final result: victory.
Appendix F Case Study Details
This appendix provides a detailed summary of the reasons for failure in each simulation conducted during the incident response game. Each point highlights the critical missteps and oversights that led to unsuccessful outcomes, offering insights into how different strategies impacted the team’s effectiveness.
  1. 1.
    Homogeneous centralized structure (seed 7): The team relied too heavily on standard, high-modifier procedures without considering nuanced indicators specific to the breach scenario. This approach resulted in a lack of adaptation to the complexities of insider threats.
  2. 2.
    Heterogeneous centralized structure (seed 1): The defenders neglected network-level indicators early in the investigation, focusing on endpoint procedures that did not align well with the attacker’s lateral movement and command and control activities. This oversight delayed detection of critical attack stages.
  3. 3.
    Heterogeneous centralized structure (seed 4): Delays in prioritizing behavior-based analytics meant the team failed to detect key indicators of insider activity, such as abnormal login patterns. This led to missed opportunities to identify the breach early.
  4. 4.
    Heterogeneous centralized structure (seed 7): The team repeatedly chose high-modifier procedures rather than aligning their strategy with the breach’s specific entry points, such as a compromised third-party partner. This mismatch resulted in an inability to address the breach’s unique characteristics.
  5. 5.
    Homogeneous decentralized structure (seed 5): Ineffective log analysis during the initial investigation caused the team to overlook early indications of credential-based attacks. This failure delayed the detection of anomalous access patterns tied to compromised credentials.
  6. 6.
    Heterogeneous decentralized structure (seed 1): By focusing narrowly on endpoint analysis, the team failed to scrutinize network-level anomalies that could have revealed attacker activity. This misstep allowed command and control communications to go undetected.
  7. 7.
    Heterogeneous decentralized structure (seed 7): The defenders repeatedly relied on established procedures with high modifiers but did not adapt their approach to the breach’s specifics. This rigid strategy hindered their ability to uncover critical details about the attacker’s movements.
  8. 8.
    Heterogeneous decentralized structure (seed 13): Insufficient focus on user behavior analytics delayed the detection of password spraying activities and insider login anomalies. The team missed an early opportunity to identify compromised credentials.
  9. 9.
    Homogeneous hybrid structure (seed 1): The failure to prioritize effective log and behavior analytics misaligned the team’s efforts with the breach’s attack pattern. This resulted in a lack of actionable insights during critical early turns.
  10. 10.
    Homogeneous hybrid structure (seed 7): The team failed to effectively leverage memory analysis, an established procedure with a high modifier, which could have uncovered critical stages of the attack. This misstep allowed stealthy malicious activity to persist undetected.
  11. 11.
    Homogeneous hybrid structure (seed 13): Initial efforts focusing on endpoint analysis did not adequately capture broader network patterns associated with lateral movement and command and control communications. This limited the team’s visibility into the attack.
  12. 12.
    Heterogeneous hybrid structure (seed 13): The defenders misprioritized their procedures, neglecting early threat vectors such as behavior analytics that could have exposed the insider threat. This oversight delayed the identification of suspicious login patterns and credential misuse.